# A Real-time Multimodal Transformer Neural Network-powered Wildfire Forecasting System

### v1

@@ 2.4 Data Uncertainties
We made an assumption when modeling it, we assume that all the data provided by the source are accurate enough including the images from Google Earth Engine. However, the reality should not be this case. In general, outliers may not affect the machine learning training results, however, some data features that have significantly large error may affect the training results. For instance, in the first part of the data, the weather location is from the nearest NOAA station that recorded weather information, some stations may have far away from the location, thus, some of the weather information may not be accurate. Moreover, the date and time when the fire was discovered may not be accurate. Wildfire is a complex process and its occurrence is the product of several interrelated factors, including ignition source, fuel composition, weather, and topography[25]. When the fire start, especially for those natural caused, it may take a while for people to find them if the fire is far away from humanity. Under this circumstance, the causes may also hard to find out, therefore, some causes may also not be accurate. As we mention in Section 2.2, the qualities of the images are varied due to the technologies. We also made an assumption that the area before wildfire occurred may not change a lot during that period. The images are stitching from Google Earth Engine, and all images are not consequently captured by satellites at the same time. Some images taken even have several years difference, for example, last image for a certain location may be capture in two years ago. Furthermore, the season when the images are being taken may also affect the results. Based on the observation of the images, we found that satellite images do not provide a normal qualities of some locations such as some place in Alaska or some place that hardly have people, which means that the image is in a quite low pixels, and it only appear with less color spectrum without showing the exact terrain. Therefore, the image data still need to be improved. Moreover, the data used in this work only contain the data of the wildfire that had occurred inside the United States. Thus, based these data, the prediction of the wildfire occurrence outside the U.S. may not be accurate.
@@ 2.5 Data and Code Availability
It is also sustainable to feed more data to obtain better results. All data could be find on official website and imagery data is also available on 1. README File. The codes, and visualization are all available on Colab. See 1. README File for the ReadMe and instructions.
@@ 3.1 Statistical Settings
For this data, we would like to predict whether a location is at risk of wildfire if it satisfies a certain natural conditions, mathematically, this is a conditional probability such that
@@
In our data, we are given that all the entries are the wildfires that have been actually occurred and the probabilities of natural causes, it can be presented as,
$\mathbb{P}(natural\ causes|{\color[rgb]{0,0,0}wildfire,\ other\ conditions})$ $=\frac{\mathbb{P}(wildfire,natural\ causes{\color[rgb]{0,0,0},\ other\ conditions})}{\mathbb{P}(wildfire{\color[rgb]{0,0,0},\ other\ conditions})},$ (2)
Finally, we can obtain Eq. (1) by using Bayes’ Theorem,
@@
When we are modeling for $\mathbb{P}(natural\ causes|wildfire{\color[rgb]{0,0,0},\ other\ conditions})$, we assume that the location has the wildfire and try to identify the probability of its reason. Among these probabilities, the probability, $\mathbb{P}(natural\ causes{\color[rgb]{0,0,0},\ other\ conditions})$
@@
and $\mathbb{P}(other\ conditions)$ shown in Eq. (3), can be predicted by other sources such as weather forecasts. Then based on the prediction of weather forecast, we are able to obtain the probabilities of natural causes(i.e. the probabilities of storm weather in a certain area). For the $\mathbb{P}(wildfire|{\color[rgb]{0,0,0}other\ conditions})$, we can refer from Probability of Ignition[26] based on NWCG. Instead directly predict the conditional probability, our model will focus on predicting the natural causes given occurrence of wildfire, for instance, the lightening caused wildfire, which is directly in the data. Once we know three of them, we cal have all probabilities. Furthermore, based on Eq. (3), $\mathbb{P}(wildfire|natural\ causes,\ other\ conditions)$ is proportional to
$\mathbb{P}(natural\ causes|wildfire,\ other\ conditions),$
@@
Thus if one of the probability increases the other one will act the same, and we can choose one to represent the other. Figure 7 summarizes the outline and data flow for the proposed model, in the previous section we collected the weather, location, date, vegetation, and image data based on the given location and date. For the information data without images, that is the features with locations, weather and vegetation categories, we process the data directly to a neural network[27]. Then this model without images will predict probabilities as outputs. In order to check the correctness of the model, we defined a classification threshold, $0.5$. Thus, if the probability is less than the threshold, we categorized it as false otherwise 1. With this metrics, we are able to understand accuracy of our models. For the data with images, inspired by a branches of CNN[28], ResNet[29] and transformer[30], we designed a feature extraction model which consists of a light version of ResNet with a modified ViT[31] with image features to exaggerate the features that it can find out but humanity cannot see by eyes and with as much low computation cost as possible, then output some new features to predict the risk of a certain location will have wildfire. Furthermore, in order to have a stable results, we first find the mean of two probabilities and then predict the probabilities in Eq. (2). In order to quantify the probabilities, we used confusion matrix to determine whether the results from model predicted are accurate or not. As we can see from Figure 7, all the features in information data and image data are based on the location and weather conditions. Then, with some feature selections and designed model to extract and enlarge the image features, we can predict the probabilities such as
@@
For a given location and weather information, the transformer machine learning model can obtain the data from the sources and calculate the probabilities. If the probabilities are larger than the threshold that we set, we can then say that for a given the nature cause occurs the fire should occur under the circumstance of weather and vegetation/terrain conditions, otherwise, the fire will not occur.
@@ 3.2 Baseline Model
In the first phase of training, the baseline model that we used is a deep neural network[32], we only used the information data to illustrate how new features would affect the results. We used several different combinations of features to see the different results, which include temperature plus wind, temperature plus wind plus humidity plus precipitation, temperature plus wind plus humidity plus precipitation plus hot-encoded vegetation. Firstly, we use only geo-referenced records which contains the weather only. We used both classification method and regression method to predict the results, since in our case we would like to have a probability and regression provided a better and stable results, thus, we used regression for now, and we shall discuss the results in the next section. For this neural network architecture, we designed 5 fully connected hidden layers, each of layer can be treated as a block. As we can see from Figure 8, after the input layer, each of these is consist of two or three parts, hidden layer, batch normalization[33] layer, and the activation function is rectified linear unit (ReLU)[34], while the last layer is another fully connect layer. Inside this model, the total amount of variables is $49705$.
@@
From the batch normalization algorithm, we can see that through a linear combinations, the mean and standard deviation of each input for the block will be within a certain range, so that the next layer do not need to satisfy the changes and increase the efficiency of the model[35]. Moreover, the model will depend less on hyper-parameter tuning so that it allows a little larger learning rate. It will also shrink the internal covariance shift[33]. Since we do not using batches in this model the batch norm layer will simply normalization the entire input matrices. With the same architecture shown in Figure 8, we appended vegetation categories and trained model again. Inside the model, as shown in Table 2, the loss function we use is the mean square error (MSE), which is a wildly used loss function (Cross Entropy loss[36] is also available in the code for classification).
@@
where $n$ is total amount of data, $y_{i}$ is the $i$th true label, and $\hat{y}_{i}$ is the $i$th predicted label by model. This metrics provides a intuitive difference between the true values and the predict values. Even though, MSE is usually not recommended for classification, we tried all other loss functions, and the model tend to predict all as false. Another way is treating the model as classification problem. Thus in the model, the loss function can be CrossEntropy loss[37], which requires one-hot encode the label and change the layer in baseline model to SoftMax[38]. We also defined a new loss function called true rate loss which will calculate both sensitivity and specificity, then maximize it. For finding gradient, we applied Adam optimizer[39].
@@
Training this model does not necessarily need a local GPU and large RAMs. It would only take about several seconds for a 100 epochs without mini-batch by using a standard NVIDIA T4 Tensor Core GPU with 16G RAM on Google Colabotory[40], where the codes and pre-trained models are also available on given Colab. For a single CPU training without mini-batch method, it will also only take less than $30$ seconds.
3.3 Wildfire Imageries Information Net (WIIN)
Besides the information data, for the image data we collected, we design a Wildfire Image Information Net, which can extracted the features from the images based on our purpose. It is a combination of a ResNet[29] liked structure and a transformer with attention mechanism liked structure. Initially, a transformer is used in translation for natural language process, thus, the self-attention mechanism focuses more about the locations of a words and its score in a certain sentence, which is an advantages that we could take. Therefore, when applying to images, a traditional vision transformer focuses on classification or segmentation. It represents the information by a token features, which is also known as class features, however, our model is not doing a classification only using images, therefore, we also need to focus more on what has been ”told” by the images. To make this work for the abandoned image pixels from the traditional vision transformer, we not only apply the self-attention mechanism, but also add a convolutional block to process the image pixels and through a multilayer perceptron so we can extract image features from it as shown in Figure 9. The reason a ResNet is applied before a transformer[30] is because usually, a transformer requires a large amount of data[41]. However, in our case instead of a large amount of data like GPTs[42] are using, we have limited data size, after undersampling method is applied we only have about $10000$ data for training. Thus, in order for the transformer can have a better performances from the image, we enlarge the features by using ResNet. ResNet have CNN[28] structures, it can extract the feature maps from the original image[43], which would make it easier for a transformer to process the feature maps instead of the original images. Moreover, based on its structures, a transformer with attention mechanism will act as a encoder to analyze these images. After embedding an image it can assign the scores and find the most possible patches that related to the target. As shown in Figure 9, we split the output features into two parts, token feature and image feature. Both features will append to the information data as new features and feed into a deep neural network for training, thus, with the original 21 features, there are totally 23 features when using the combined model.
@@ 3.3.1 Residual Network
At this stage, we will extract feature maps from original captured image so that we can both maintain the performance of the model and save computation power. As we can see from Figure 9, the mini-batched images with size of $(100,100)$ will first go through a convolutional block, which includes a convolutional layer, a batch normalization layer, an activation layer, and a maxpooling layer, as shown in left of Figure 10 (a). The convolutional layers are used to extracted features from any previous procedures and it will output a new feature map of the previous stage of image or feature map [43]. Mathematically, for a grey-scaled image $I\in\mathbb{R}^{W\times H\times 1}$, where W stands for width of the image and H is the height of the image, the ij-th entry of output feature map by using the kernel $K\in\mathbb{R}^{p\times q}$ can be represented as [44]
@@
In this stage, the convolutional layer has a kernel size of 5, stride of 2, and padding of 3. Taking Figure 10 (b) as an example, the blackboxes with 0s are the padding area and the rest are pixels of the original image. The $5\times 5$ yellow and green shaded areas will multiply weighted kernels that will be trained in different time with each time the kernel will move 2 boxes since the stride is set to 2. Similar to convolutional layer, the maxpooling layer will keep main features and reduce parameters in the model, it removes all the value except for the maximum value in each filter process [43]. It will then obtain the maximum numbers within each kernel instead of multiply the trained weights. Therefore, after the convolutional block the image size will become $(51,51)$ as unnecessary features would be removed.
@@
Then, the convoluted images will be fed into two residual block with 3 layers and 2 layer separately. The residual blocks are mainly two convolutional blocks without maxpooling layer but add with downsampled, if necessary, initial inputs as residual. As shown in right of Figure 10 (a), for a single residual block, the new residual $F(x)$ will add back to initial input value $x$ to create a new output $F(x)+x$. With the help of adding residuals to the output, ResNet can provide a good feature map from original image without losing too much information and prevent gradient vanish or explode[45, 29]. Even we have a small size of image, it can still work well. After two residual blocks, the image size reduce to $(13,13)$, then continue to feed the processed images into the last maxpooling layer. Mathematically, let $F(\cdot)=BN(Conv(\cdot))$
@@ 3.3.2 Wildfire Images Transformer (WIT)
In the transformer stage, similar to a vision transformer (ViT) (See: [31]), the output image from previous stage will be fed into a convoluted embedded patch layer so that the image features will be flatten, as shown in Figure 11, meanwhile, an learnable feature token will be appended to the newly flatten features, as shown in left of Figure 12. Mathematically, this block with output shown in Figure 11 can represented as
@@
For instance, for an image in our output of ResNet, $x_{i}\in\mathbb{R}^{1\times 10\times 10}$, with the patch size of 2, the flattened vector should have $\frac{H\times W}{P^{2}}=\frac{100}{4}=25$ patches concatenate an additional token vector and plus positional encoding. The output dimension will become a $x\in\mathbb{R}^{(25+1)\times 4}$ features. Our model applied a convolution method to the image since our patch size is small, only $4$, with the output channel $64$ thus the output dimension of embedded patch layer is $(26,64)$.
@@
Then, the data will have two transformer blocks and each of them mainly have a multi-head attention layer and a multilayer perception, as shown in the right of Figure 12. In multi-head attention layer, the flattened input feature vectors with appending tokens will apply linear projections by multiply three different weight, $w_{i}\in\mathbb{R}^{P^{2}C\times\frac{P^{2}C}{n\_head}},i\in\{q,k,v\}$ to three new vectors $Q(query)$, $K(key)$, and $V(value)$, for instance $w_{q}x=Q$. Thus, in this case, Query, key, and value vectors should have shape $(b,26,8,8)$ as the hidden dimension is $64(8\times 8)$. The $n\_head$ stands for number of heads defined for multi-head attention. Vaswani et al.[30]. define the output of the attention layer as concatenating the various heads and then multiplying by a matrix $w^{o}$.
@@
where $w^{o}=\begin{bmatrix}w^{o}_{1}\\ \vdots\\ w^{o}_{h}\end{bmatrix}$ As the results are concatenated, the dimension is $\frac{P^{2}C}{n\_head}\times n\_head=P^{2}C$. Thus, at the end of this layer the dimension is still $\mathbb{R}^{P^{2}C\times P^{2}C}$. Once $Q,K,V$ are obtained, the score could be computed by $softmax(QK^{T})$, and logit scaled by $\frac{1}{\sqrt{d_{qkv}}}$ for scaled dot-product score as follows,
@@
and a dropout layer to prevent overfitting. In the last, linear layer the output dimension will be changed back to its origin. Thus the dimension of the vectors has no changes. Instead of output a token class only like ViT, our model will also output image features which is can be treat as summarized features of an image. Thus, the output of the wildfire image information network stage, we will have extra features for appending to the neural network model.

### v2

@@ 2.4 Data Uncertainties
In data processing and characterization, we assume that all the data provided by the source is accurate enough including the images from Google Earth Engine. It contains less epistemic uncertainty, if any. In reality, however, this may not be the case. In general, outliers may not affect the machine learning training results, however, some data features that have significantly large error may affect the training results. For instance, in the first part of the data, the weather condition information at a specific location may be recorded from the nearest National Weather Service Station (NWS) or the nearest National Oceanic and Atmospheric Administration (NOAA) station, while some stations may be far away from the location, Thus, some of the weather information may not be accurate. Moreover, the date and time when the fire was discovered may also not be accurate. Wildfire is a complex process and its occurrence is the product of several interrelated factors, including ignition source, fuel composition, weather, and terrain topography feature [25]. When a fire starts, especially for those nature-caused fires, it may take a while for people to find or discover them, if the fire is far away from society. Under those circumstances, the causes of the fires may also be hard to find out, therefore, some causes may also not be accurate. As mentioned in Section 2.2, the qualities of the images are varied due to the technologies. Another assumption is also made that the area before a wildfire occurred may not change a lot during that period. The images are stitched from Google Earth Engine, and all pieces of an image are not consequently captured by satellites at the same time. Some images taken even have several years difference, for example, the last image for a certain location may be captured two years ago. Furthermore, the season when the images are being taken may also affect the results. Based on the observation of the images, satellite images do not provide normal qualities of some locations such as some place in Alaska or some places that hardly have people, which means that the image has quite low resolution, and some images only appeared with less color spectrum without showing the exact terrain. Therefore, the quality of image data still needs to improve. Moreover, the data used in this work only contain the data of the wildfire that had occurred inside the United States. Thus, based these data, the prediction of the wildfire occurrence outside the U.S. may not be accurate.
@@ 2.5 Data and Code Availability
It is both preferable and sustainable to feed more data to train the AI model to obtain better results. All training data can be found on the official websites, and imagery data is also available on 1. README File. The computer codes for generating visualization are available on Colab. See 1. README File for the ReadMe and instructions.
@@ 3.1 Statistical Settings
The technical objective of this work is to predict whether a location is at risk of wildfire occurrence. The prediction is characterized by a set of conditions, natural or man-made. Mathematically, this prediction may be expressed in terms of conditional probability defined as follows,
@@
In our data format, all the data entries are related to the wildfires that have been actually occurred. By virtue of probability chain rule, the condition probability of a wildfire occurrent due to natural causes, or other conditions can be expressed as follows,
$\mathbb{P}(natural\ causes|{\color[rgb]{0,0,0}wildfire,\ other\ conditions})$ (2)
$=$ $\frac{\mathbb{P}(wildfire,natural\ causes{\color[rgb]{0,0,0},\ other\ conditions})}{\mathbb{P}(wildfire{\color[rgb]{0,0,0},\ other\ conditions})},$
Finally, we can reformulate Eq. (1) by using Bayes’ Theorem,
@@
When we are modeling for $\mathbb{P}(natural\ causes|wildfire{\color[rgb]{0,0,0},\ other\ condiions})$, we assume that the location has had a wildfire and try to identify the conditional probability of its cause. Among these probabilities, the probabilities, $\mathbb{P}(natural\ causes{\color[rgb]{0,0,0},\ other\ conditions})$
@@
and $\mathbb{P}(other\ conditions)$ shown in Eq. (3), can be predicted by other sources such as weather forecasts. Then based on the prediction of weather forecast, we can obtain the probabilities due to natural causes, i.e. the probabilities of storm weather in a certain area. For the $\mathbb{P}(wildfire|{\color[rgb]{0,0,0}other\ conditions})$, we can refer from Probability of Ignition[26] provided by NWCG. Instead of directly predicting the conditional probability, our model will focus on the prediction of the likelihood probability of the natural causes given the occurrence of wildfire, for instance, the lightening caused wildfire, which can be directly found in the data. Thus, once we know three of them, we can have all the probability needed to make a posterior prediction. Furthermore, based on Eq. (3), we have
$\mathbb{P}(wildfire|natural\ causes,\ other\ conditions)$
$\propto\mathbb{P}(natural\ causes|wildfire,\ other\ conditions),$
@@
Thus if one of the probabilities increases the other one will act the same, and we can choose one to represent the other. Figure 7 is a flowchart that outlines data flow for the proposed model discussed in the previous Section. For each past wildfire record, we collected the weather, location, date, vegetation, and image data based on the given location and date. For the information data without images, which is the features associated with locations, weather, and vegetation categories, we process the data directly to a neural network[27]. Then our model without images can also provide predicted probabilities as outputs. To validate the correctness of the proposed model, we defined a classification threshold, $0.5$. Thus, if the probability is less than the threshold, we categorize it as false otherwise 1. With these metrics, we can assess the accuracy of the proposed model. For the data with images, inspired by architectures of CNN[28], ResNet[29], and Transformer[30], we designed and implemented a feature extraction model that consists of a light version of ResNet with a modified ViT[31] with image features to exaggerate the features that it can discern but humans cannot see by naked eyes and with as much low computation cost as possible. By doing so, we can then output some new features to predict the risks of wildfire at certain locations. Moreover, to have a stable results, the model will first find the mean of two probabilities and then predict the probabilities in Eq. (2). To quantify the probabilities, the confusion matrix is used to determine whether the results predicted by the model are accurate or not. As shown from Figure 7, all the features in information data and image data are based on the location and weather conditions. Hence, with some feature selections, a designed model is used to extract and enlarge the image features, so that we can predict the probabilities such as
@@
For the given location and weather information, the transformer machine learning model can obtain the data from the sources and calculate the probabilities. If the probabilities are larger than the threshold that we set, we can then say that if a given natural cause occurs the fire should occur under the circumstance of the weather and vegetation/terrain conditions. Otherwise, the fire will not occur.
@@ 3.2 Baseline Model
In the first phase of the training, the baseline model used is a deep learning neural network[32], in which only the information data is used to illustrate how new features would affect the results. We used several different combinations of features to observe the result or output variations, which include temperature plus wind, temperature plus wind plus humidity plus precipitation, temperature plus wind plus humidity plus precipitation plus hot-encoded vegetation, etc. First, only geo-referenced records would be fed into the model. Both the classification method and regression method are used to predict the probability results. Since in this case, one would like to have a probability result and the regression strategy can provide better and more stable results, thus, regression strategy is used for the task, and the results shall be discussed in the next section. For this neural network architecture, the model has five fully connected hidden layers, and each layer can be treated as a block. As shown in Figure 8, after the input layer, each of these consists of two or three parts, hidden layer, batch normalization layer [33]. Here, the activation function is rectified linear unit (ReLU)[34], while the last layer is another fully connected layer. Inside this model, the total amount of variables is $49705$.
@@
From the batch normalization algorithm, through a linear combination, the mean and standard deviation of each input for the block will be within a certain range, so that the next layer does not need to satisfy the changes and increase the efficiency of the model[35]. Moreover, the model will depend less on hyper-parameter tuning so that it allows a little larger learning rate. It will also shrink the internal covariance shift[33]. From the batch normalization algorithm, we can see that through a linear combinations, the mean and standard deviation of each input for the block will be within a certain range, so that the next layer does not need to be adjusted with possible changes and increase the efficiency of the model[35]. Moreover, the model will depend less on hyper-parameter tuning so that it allows a little larger learning rate. It will also shrink the internal covariance shift[33]. Since we do not use batches in this model, the batch norm layer will be simply the normalization of the entire input matrices. With the same architecture shown in Figure 8, we appended vegetation categories and trained the model again. Inside the model, as shown in Table 2, the loss function we use is the mean square error (MSE), which is a wildly used loss function, and the Cross Entropy loss[36] is also available in the code for classification.
@@
where $n$ is total amount of data, $y_{i}$ is the $i$th true label, and $\hat{y}_{i}$ is the $i$th predicted label by model. This metric provides an intuitive difference between the true values and the predicted values. Another way is treating the model as a classification problem. Thus in the model, the loss function can be the Cross-Entropy loss function [37], which requires one-hot encoding of the label and changing the layer in the baseline model to SoftMax[38]. A new loss function called true rate loss will calculate both sensitivity and specificity are defined, and then maximized it. For finding the gradient, the Adam optimizer[39] is used.
@@
Training this model does not necessarily need a local GPU and large RAMs. It would only take about several seconds for 100 epochs without a mini-batch by using a standard NVIDIA T4 Tensor Core GPU with 16G RAM on Google Colabotory[40], where the codes and pre-trained models are also available on given Colab. For a single CPU training without a mini-batch method, it will also only take less than $30$ seconds.
3.3 Wildfire Imagery Information Net (WIIN)
Besides the information data, for the imagery data collected, we designed a Wildfire Image Information Net, which can extract the features from the images according to our purpose. It is a combination of a deep learning architecture of ResNet[29] type and a transformer with an attention-based architecture. Initially, a transformer is used in translation for the natural language process, thus, the self-attention mechanism focuses more on the locations of a word and its score in a given sentence, which is beneficial for understanding complex language structures. When applying the transformer to images, a conventional vision transformer focuses on classification or segmentation. It represents the information by token features, which are also known as class features. However, our model is not concerned with classification but only interested in the information embedded in images, Therefore, the proposed model focuses more on what has been “told” by an image. To make this work for the abandoned image pixels from the conventional vision transformer, the model not only applies the self-attention mechanism but also adds a convolutional block to process the image pixels through a multilayer perceptron so that the model can extract image features from it as shown in Figure 9. The reason why the ResNet is applied before the transformer[30] is because usually, a transformer requires a large amount of data[41]. However, in our case, instead of using a large amount of data like GPTs[42] are using, we have limited training data size, after the undersampling method is applied, we only have about $10000$ data for training. Thus, in order for the transformer can have better performance in image processing, we enlarge the features by using ResNet. ResNet has CNN-like[28] architecture, it can extract the feature maps from the original image[43], which would make it easier for a transformer to process the feature maps instead of the original images. Moreover, based on its architecture, a transformer with attention mechanisms will act as an encoder to analyze these images. After embedding an image, it can assign the scores and find the most possible patches that are related to the target. As shown in Figure 9, we split the output features into two parts: the token feature and the image feature. Both features will append to the information data as new features and feed into a deep neural network for training. Thus, with the original 21 features, there are a total of 23 features when using the combined model.
@@ 3.3.1 Residual Network
At this stage, feature maps from the original captured image are extracted so that the model can both maintain its performance and save computation power. As shown from Figure 9, the mini-batched images with the size of $(100,100)$ will first go through a convolutional block, which includes a convolutional layer, a batch normalization layer, an activation layer, and a max-pooling layer, as shown on the left of Figure 10 (a). The convolutional layers are used to extract features from any previous procedures, and it will output a new feature map of the previous stage of an image or feature map [43]. Mathematically, for a grey-scaled image $I\in\mathbb{R}^{W\times H\times 1}$, where W stands for the width of the image, and H is the height of the image, the ij-th entry of output feature map by using the kernel $K\in\mathbb{R}^{p\times q}$ can be represented as [44]
@@
In this stage, the convolutional layer has a kernel size of 5, stride of 2, and padding of 3. Taking Figure 10 (b) as an example, the blackboxes with 0s are the padding area, and the rest are pixels of the original image. The $5\times 5$ yellow and green shaded areas will multiply weighted kernels that will be trained at different times, and each time, the kernel will move two boxes since the stride is set to 2. Similar to the convolutional layer, the maxpooling layer will keep the main features and reduce parameters in the model, it removes all the values except for the maximum value in each filter process [43]. It will then obtain the maximum numbers within each kernel instead of multiplying the trained weights. Therefore, after the convolutional block, the image size will become $(51,51)$ as unnecessary features would be removed.
@@
Then, the convoluted images will be fed into two residual blocks with three and two layers separately. The residual blocks are mainly two convolutional blocks without the max-pooling layer, added with downsampled initial inputs as residuals if necessary. As shown in right of Figure 10 (a), for a single residual block, the new residual $F(x)$ will add back to initial input value $x$ to create a new output $F(x)+x$. With the help of adding residuals to the output, ResNet can provide a good feature map from the original image without losing too much information and prevent the gradient vanish or exploding [45, 29]. Even if the fed data has a small image size, it works well. After two residual blocks, the image size reduces to $(13,13)$, and we then feed the processed images into the last max-pooling layer. Mathematically, let $F(\cdot)=BN(Conv(\cdot))$ Similar to the convolutional layer, the max-pooling layer will keep the main features and reduce parameters in the model, it removes all the values except for the maximum value in each filter process [43]. It will then obtain the maximum numbers within each kernel instead of multiplying the trained weights. Therefore, after the convolutional block, the image size will become $(51,51)$ as unnecessary features would be removed. Then, the convoluted images will be fed into two residual blocks with three and two layers separately. The residual blocks are mainly two convolutional blocks without the max-pooling layer, added with downsampled initial inputs as residuals if necessary. As shown in right of Figure 10 (a), for a single residual block, the new residual $F(x)$ will add back to initial input value $x$ to create a new output $F(x)+x$. With the help of adding residuals to the output, ResNet can provide a good feature map from the original image without losing too much information and prevent the gradient vanish or exploding [45, 29]. Even if the fed data has a small image size, it works well. After two residual blocks, the image size reduces to $(13,13)$, and we then feed the processed images into the last max-pooling layer. Mathematically, let $F(\cdot)=BN(Conv(\cdot))$
@@ 3.3.2 Wildfire Images Transformer (WIT)
In the transformer stage, similar to a vision transformer (ViT) (See: [31]), the output image from the previous stage will be fed into a convoluted embedded patch layer so that the image features will be flattened, as shown in Figure 11, meanwhile, a learnable feature token will be appended to the newly flattened features, as shown in the left of Figure 12. Mathematically, this block with the output shown in Figure 11 can represented as
@@
For instance, for an image in the output of ResNet, $x_{i}\in\mathbb{R}^{1\times 10\times 10}$, with the patch size of 2, the flattened vector should have $\frac{H\times W}{P^{2}}=\frac{100}{4}=25$ patches concatenate an additional token vector and positional encoding. The output dimension will become a $x\in\mathbb{R}^{(25+1)\times 4}$ features. Our model applied a convolution method to the image since our patch size is small, only $4$, with the output channel $64$. Thus, the output dimension of the embedded patch layer is $(26,64)$.
@@
Then, the data will have two transformer blocks, each of which will have a multi-head attention layer and a multilayer perception, as shown in the right of Figure 12. In the multi-head attention layer, the flattened input feature vectors with appending tokens will apply linear projections by multiplying three different weight, $w_{i}\in\mathbb{R}^{P^{2}C\times\frac{P^{2}C}{n\_head}},i\in\{q,k,v\}$ to three new vectors $Q(query)$, $K(key)$, and $V(value)$, for instance $w_{q}x=Q$. Thus, in this case, Query, key, and value vectors should have shape $(b,26,8,8)$ as the hidden dimension is $64(8\times 8)$. The $n\_head$ stands for the number of heads defined for multi-head attention. Vaswani et al.[30]. define the output of the attention layer as concatenating the various heads and then multiplying by a matrix $w^{o}$.
@@
where $w^{o}=\begin{bmatrix}w^{o}_{1}\\ \vdots\\ w^{o}_{h}\end{bmatrix}$ As the results are concatenated, the dimension is $\frac{P^{2}C}{n\_head}\times n\_head=P^{2}C$. Thus, at the end of this layer, the dimension is still $\mathbb{R}^{P^{2}C\times P^{2}C}$. Once $Q,K,V$ are obtained, the score could be computed by $softmax(QK^{T})$, and logit are scaled by $\frac{1}{\sqrt{d_{qkv}}}$ for the scaled dot-product score as follows,
@@
and a dropout layer to prevent overfitting. In the last, linear layer the output dimension will be changed back to its origin. Thus the dimension of the vectors has no changes. Instead of outputting a token class only like ViT, our model will also output image features, which can be treated as summarized features of an image. Thus, at the output stage of the wildfire image information network, we will have extra features for appending to the neural network model.